# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v1

@@ C.3 Proof of Theorem 3
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star}\|_{L^{2}(\mu)}\leq\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}+C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$
which is the slab-wise version of (71) (with $\rho_{*}$ in place of a global $C_{1}$). ∎
Remark 4 (Relation to prior PGDPO slab analyses).
The closure step above uses the same slab-wise small-gain philosophy as in huh2025breaking: short-time parabolic smoothing yields a contraction on each time slab, and the global bound follows by concatenation. The key difference here is that the contraction is applied to the $q$-aggregated projection map $T(\pi)=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$, hence the additional use of the algebraic projection stability (Proposition 1) for the estimator $\widehat{\pi}^{\mathrm{agg,mix}}$.
Appendix D Implementation details for Section 3
@@
D.1 Stage 1 (PG–DPO) template for Section 3.1
@@
D.2 Stage 2 (P–PGDPO projection; mixed-moment $q$-aggregation) template for Section 3.2
@@
D.3 Coupling I: residual/control-variate projection (Section 3.3.1)
@@
Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; query state(s) $z=(t,x,y)\sim\mu$; and stage 2 projection ingredients $(\widehat{A}_{t}(z),\widehat{G}_{t}^{\mathrm{mix}}(z))$ constructed as in Section D.2.
@@
D.4 Coupling II: interactive distillation (Section 3.3.2)
@@
      Stage 1 update (PG–DPO step). Perform one PG–DPO update step on $J(\varphi)$ as in Section D.1.
@@
        Optionally filter states using diagnostics (Section D.5) and update the buffer:
@@
D.5 Engineering notes and stabilizers
@@
Appendix E Stage 2 projection diagnostics
@@
Appendix F Supplementary decomposition diagnostics for Section 5

### v2

@@ D.3 Proof of Theorem 3
$\|\widehat{\pi}_{\mathrm{agg,mix}}-\pi^{\star}_{\mathrm{blind}}\|_{L^{2}(\mu)}\leq\|\widehat{\pi}_{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\|_{L^{2}(\mu)}+\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}}+\frac{\rho^{\ast}\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm}},$
which is the claimed slab-wise residual-based $\theta$-blind policy-gap bound. ∎
Appendix E Implementation details for Section 3
@@
E.1 Stage 1 (PG–DPO) template for Section 3.1
@@
E.2 Stage 2 (P–PGDPO projection; mixed-moment $q$-aggregation) template for Section 3.2
@@
E.3 Coupling I: residual/control-variate projection (Section F.1)
@@
Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; query state(s) $z=(t,x,y)\sim\mu$; and stage 2 projection ingredients $(\widehat{A}_{t}(z),\widehat{G}_{t}^{\mathrm{mix}}(z))$ constructed as in Section E.2.
@@
E.4 Coupling II: interactive distillation (Section F.2)
@@
      Stage 1 update (PG–DPO step). Perform one PG–DPO update step on $J(\varphi)$ as in Section E.1.
@@
        Optionally filter states using diagnostics (Section E.5) and update the buffer:
@@
E.5 Engineering notes and stabilizers
@@
Appendix F Coupling stage 1 and stage 2: residual projection and interactive distillation
We keep the ex–ante objective (48) and the $\theta$-blind deployability constraint throughout. Stage 2 is not a separate optimization problem: it is a (warm-started) post-processing map that reuses the current stage 1 policy as a warm-up control, estimates the projection ingredients under this warm-up policy, and then applies the $q$-aggregated Pontryagin projection as a deterministic transformation of the policy. Concretely, stage 2 takes $\pi^{\mathrm{warm}}$ as input, runs BPTT/Monte Carlo under frozen $\theta\sim q$ to estimate the adjoint blocks required for the Hamiltonian gradient, aggregates them across $\theta$, and outputs a single deployable $\theta$-blind rule $\widehat{\pi}^{\mathrm{agg,mix}}$ (Section 3.2).
This appendix records two couplings between the two stages, each with a distinct role:
  • Residual (control-variate) projection. We compute the same projected rule in a residual form around the warm-up policy. This is algebraically equivalent to the direct projection, but typically reduces Monte Carlo variance and improves numerical stability in high dimensions.
  • Interactive distillation (amortized projection). We use the projected rule as a teacher signal at intermittent refresh times. Beyond accelerating/stabilizing training, distillation amortizes the cost of projection: stage 2 can be accurate but Monte-Carlo intensive, whereas a distilled student policy approximates the projected rule with a single forward pass at stage 1 inference cost.
F.1 Control-variate (residual) form of the projected rule
Recall the mixed-moment projected rule (59). In high dimensions, Monte Carlo noise in the projection inputs can be non-negligible, and solving a linear system with $\widehat{A}_{t}$ can amplify this noise. A convenient stabilization is to compute the same projected rule in a residual (control-variate) form around the warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$.
and the aggregated residual (the quantity we actually solve against)
Whenever $\widehat{A}_{t}(t,x,y)$ is (numerically) invertible, the projected rule admits the identity
which is an algebraic rewriting of (59) (hence it does not change the target).
Practical value (variance reduction and stability). When the warm-up policy is already close to a projected fixed point on the working domain, the residual $\widehat{r}_{\mathrm{FOC}}$ tends to be small. Moreover, the ingredients entering $\widehat{A}_{t}\pi_{\varphi^{\mathrm{warm}}}$ and $\widehat{G}_{t}^{\mathrm{mix}}$ are computed from the same Monte Carlo pool and often partially cancel, improving concentration of $\widehat{r}_{\mathrm{FOC}}$. In implementation, we combine (104) with the inversion stabilizers of Appendix E.5 (e.g. ridge regularization and diagnostic-based fallback to $\pi_{\varphi^{\mathrm{warm}}}$ on unreliable states).
F.2 Interactive distillation: projection-guided training and amortized deployment
Let $\pi_{\varphi}$ be the trainable stage 1 policy network. At intermittent refresh times, we freeze a lagged copy $\pi_{\varphi^{-}}$ and run stage 2 under $\pi_{\varphi^{-}}$ to construct a $q$-aggregated projected teacher. This coupling serves two purposes:
  • Training aid. Projection-guided targets can stabilize and accelerate stage 1 optimization, especially once the warm-up policy enters a locally stable regime.
  • Amortization. Stage 2 projection can be accurate but Monte-Carlo intensive; distillation compresses it into a single forward pass for deployment.
Teacher definition (computed by stage 2, treated as fixed).
In residual form (104), the teacher is the $\theta$-blind map
where $\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}:=\widehat{A}_{t}^{\varphi^{-}}\pi_{\varphi^{-}}+\widehat{G}_{t}^{\mathrm{mix},\varphi^{-}}$ is computed using the mixed-moment $q$-aggregation under the lagged policy. In practice, the inversion in (105) is implemented with the same stabilizers as in Section 3.2 (ridge/diagnostics/residual form); for notational simplicity we keep the formal inverse.
Student objective (hybrid PG–DPO + distillation loss).
We train $\pi_{\varphi}$ by combining the original ex–ante objective with a proximity term to the teacher on the working domain:
where $\mu$ is the working-domain sampling measure and $\lambda\geq 0$ controls the strength of projection guidance. The operator $\mathrm{stopgrad}(\cdot)$ indicates that gradients are not propagated through stage 2: once computed from $\pi_{\varphi^{-}}$, the teacher is treated as fixed.
Refresh schedule, annealing, and diagnostic gating (implementation).
In practice, $\varphi^{-}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}$ are refreshed on a slower timescale than the stage 1 gradient steps. A typical and robust configuration is:
  • Two-time-scale refresh. Hold $\varphi^{-}$ fixed for $K_{\mathrm{refresh}}$ stage 1 updates of $\varphi$ under (106), then set $\varphi^{-}\leftarrow\varphi$ and recompute the teacher via stage 2.
  • Anneal $\lambda$. Start with $\lambda=0$ (pure PG–DPO) and increase $\lambda$ only after basic projection checks on the working domain indicate that the stage 2 map is reliable. Concretely, we anneal $\lambda$ toward a target value $\lambda_{\max}$ once the warm-up residual $\|\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{L^{2}(\mu)}$ and curvature diagnostics remain stable over several refresh cycles.
  • Adaptive teacher selection (gating). To avoid injecting noisy teacher targets early in training or on pathological regions of the domain, we optionally apply the teacher loss only on states where projection diagnostics certify reliability (e.g. bounded condition number of $\widehat{A}_{t}$, eigenvalue/denominator lower bounds, and moderate residual magnitude), falling back to pure PG–DPO elsewhere.
Implementation details, including the precise diagnostics and stabilizers, are deferred to Appendix E.
Deployment. After training, the deployed controller is simply the student network $\pi_{\varphi}$, which approximates the stage 2 projected rule with a single forward pass, without repeatedly running BPTT/Monte Carlo projection at inference time.
Appendix G Stage 2 projection diagnostics
@@
Appendix H Supplementary decomposition diagnostics for Section 5